OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y = 3/sin x + cos x, P = (0, 3)

OpenStudy (anonymous):

do u know what will be the slope of tangent line?

OpenStudy (anonymous):

I know that you have to differentiate the equation, then using x solve it which will result in the slope but I keep getting the wrong answer

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

x=0 is not in the Domain of y = 3/sin x + cos x !!!

OpenStudy (anonymous):

so does that mean that the answer is 0?

OpenStudy (anonymous):

or would it be y=-3x-3

OpenStudy (anonymous):

that mens y = 3/sin x + cos x is not continuous at all in x=0 and u cant find a tangent line

OpenStudy (anonymous):

or maybe im wrong @waterineyes

OpenStudy (anonymous):

the answer was -3x+3 I must've done the math wrong and made a mistake

OpenStudy (anonymous):

I have no clue about this... @mukushla Sorry.. dpalnc will clear all the doubts..

OpenStudy (anonymous):

is this your function: \(\large y=\frac{3}{sinx + cosx} \)

OpenStudy (anonymous):

if no, then @mukushla is correct... x=0 is not in the domain..

OpenStudy (anonymous):

yes it was but I submitted that one and made a silly algebraic mistake. can you help me with another one: Find an equation of the tangent line to the given curve at the specified point. 5x/x+2 at point 3,3

OpenStudy (anonymous):

the procedure is the same as the previous one... you'll need to find the derivative...

OpenStudy (anonymous):

just don't make the same algebraic mistake....

OpenStudy (anonymous):

show us what you've done and we can see where your error might be....

OpenStudy (anonymous):

at least tell us what you have for the derivative...

OpenStudy (anonymous):

i have the slope as 7/25 when I found the derivative but I'm not sure if thats right

OpenStudy (anonymous):

at x=3 your slope is not correct... what do you have for the derivative?

OpenStudy (anonymous):

-5x/(x+2)^2

OpenStudy (anonymous):

i got something else for the derivative... the denominator is correct though...

OpenStudy (anonymous):

wanna work out the derivative with me?

OpenStudy (anonymous):

is there a -5x in it?

OpenStudy (anonymous):

in the numerator*

OpenStudy (anonymous):

when simplified, no, the numerator should be some constant.

OpenStudy (anonymous):

i set it up as: x+2 (d/dx) 5x - 5x (d/dx) x+2

OpenStudy (anonymous):

the derivative of 5x is 1and so is the derivative of x+2 so its x + 2 -5x

OpenStudy (anonymous):

derivative of 5x is 5

OpenStudy (anonymous):

ok... this is just the numerator.... \(\large (x+2)[5x]'-(5x)[x+2]' \) \(\large =(x+2)\cdot 5-(5x)\cdot 1 \) \(\large =5x+10-5x \) \(\large =10\)

OpenStudy (anonymous):

thank you so much now I understand

OpenStudy (anonymous):

so you have the slope of the tangent line at x=3?

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

yes... and use that along with the point (3, 3) to find point-slope form for the equation of the tangent line... :)

OpenStudy (anonymous):

2/5x+9/5

OpenStudy (anonymous):

can you help me with another one:

OpenStudy (anonymous):

sure.. close shop on this question and post up another.... this one is making me dizzy going up and down....:|

OpenStudy (anonymous):